May you please solve these systems of equations?
Y=2x-3
And y=3x-4

Answer by sudhanshu_kmr(1152)

y = 2x -3
y = 3x-4
so, 2x -3 = 3x-4
=> x = 1
y = 2*1 -3 = -1
so, x = 1, y = -1